OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

for perpendicular line the product of their gradients is -1

OpenStudy (anonymous):

Ok so what would be the equation ?

OpenStudy (anonymous):

line 2 has a slope of 4 so line 1 which passes through (2,-3) will have a slope of -1/4 to satisfy the condition and we know that y=mx+c then we first find c(y-intercept) as below substitute y with -3,x with 2 and m with -1/4 -3=-1/4*2+c c=-3+1/2 c=-5/2 so equation will be y=-1/4x-5/2

OpenStudy (anonymous):

Thank You !

OpenStudy (anonymous):

anytime!!!!